A solution of 75% pesticide is to be mixed with a solution of 51% pesticide to form 48 liters of a 64% solution. How much of the 75% solution is needed?
Answer by mangopeeler07(462)

A solution of 75% pesticide is to be mixed with a solution of 51% pesticide to form 48 liters of a 64% solution. How much of the 75% solution is needed?

x=liters of 75%
48-x=liters of 51%
coefficient=percent

75(x)+51(48-x)=64(48)

Distribute
75x+2448-51x=3072

Combine like terms
24x+2448=3072

Subtract 2448 from both sides
24x=624

Divide both sides by 24
x=26

x=liters of 75%
48-x=liters of 51%

Answer:
liters of 75%=26
liters of 51%=22